OpenStudy (anonymous):

Don't know what i am doing at all... Help Please? Input in standard form the equation of the given line. The line that passes through (-2, 4) and is parallel to x - 2y = 6 (Standard form is Ax+By=C)

OpenStudy (anonymous):

put this x - 2y = 6 in slope-intercept form. parallel lines have the same slope so when it's in that form, we can easily see what the slope is

OpenStudy (anonymous):

y=-1/2x-3

OpenStudy (anonymous):

x - 2y = 6 -2y = -x + 6 y = 1/2x - 3

OpenStudy (anonymous):

so what's the slope?

OpenStudy (anonymous):

1/2 would be the slope...

OpenStudy (anonymous):

ok so take (-2, 4) and 1/2 and put them into point-slope form

OpenStudy (anonymous):

alrighty i got it from there, i was so confused thank you for the help (:

OpenStudy (anonymous):

no problem :] make sure you change from point-slope form to standard form!